Project

General

Profile

05 19 14

Attending: Craig, Gavin, Nate, Michael, Pavin, Dominik, Satish, Jan, Chris, Matthew, Eric,

Kanika:
— Can we update to the new ART today?
— Make sure to make a tag right before the update.
— We would prefer to wait until Thursday and make sure Neutrino storm has passed.

Craig:
— Sent around draft of FTS request — Need to iterate and submit.
— Working on updating DocDB 10129

LEM bug:
— Committed a fix, but the bug killed about 5% of calculations.
— May have a 50% chance of killing one of the cosmics that make it through pre selection.
— Should we do the reprocessing?

Eric and Nate (loose FTS):
— 1373 GENIE completed (PID and LEMSUM)
— 9500 CRY
(600 error processes)
— LEM — Chris ran a bunch of that already
— 3512 done for Cry
— No GENIE…

Dominick:
— Data is done (4 DB partition 1 data)

PCList:
— All caught up back to Jan.
— Feature request —> Get everything that is “recent”
— Satish will file a feature request.

Production has been running great, a few disclaimer:
— VM02 —> “went to sleep” a few times.
— Load balancing on the FTS (we ready to implement it even without a feature request)
—> Dominick will work on this (update Thursday?)
— Memory loads
— Database memory spikes?
— Small files —>
— Improve Reco/Sim/PID validation steps

RunNovaSam currently doesn’t support writing back histograms.

SawWise response shifts:
———————————
Monday: Dominick
Tuesday: Nate
Wednesday: Gavin,
Thursday: Eric
Friday: Adam

ND overlay:
— Update on Thursday—>